OpenStudy (deercult):

19. Solve the equation. Check for extraneous solutions. 9|9-8x| = 2x + 3

OpenStudy (will.h):

Rock is already replying so you don't need me ;)

OpenStudy (zzr0ck3r):

you need to solve two things \(9-8x=\dfrac{2x+3}{9}\) and \(9-8x=-(\dfrac{2x+3}{9})\)

OpenStudy (vuriffy):

@zzr0ck3r Aren't you able to graph it as well? Curious, I found the two values myself, wanted to see if my method worked as well with your final answer at the end.

OpenStudy (zzr0ck3r):

\(|f(x)|=a\implies f(x)=a \text{ and } f(x) =-a\)

OpenStudy (deercult):

Ohhh. That immediately explains why I keep getting it wrong, tbh

OpenStudy (zzr0ck3r):

You can graph it but it is going to be hard to tell where they intersect unless it is a nice number, and it is not.

OpenStudy (vuriffy):

Working out just fine for you, deercult?

OpenStudy (deercult):

Eh, no. I'm miserable with math, I think now I'm just messing up the order of equations, though? But in general, I doubt 0.185... is supposed to be an answer. I tried to work it out like this: 9 - 8x = (2x+3)/9 6 - 8x = 2x / 9 54 - 8x = 2x 54 = 10x And I think I'm gonna try and redo it again. Rah.

OpenStudy (vuriffy):

I don't see why you did -3 first.

OpenStudy (vuriffy):

I did my graphing method and already have both x values, but let me try this method he showed:

OpenStudy (deercult):

Oh. I got PEMDAS backwards, incredible

OpenStudy (vuriffy):

9 - 8x = (2x + 3 ) / 9 multiply by 9 81 - 72x = 2x + 3 (add 72x) 81 = 74x + 3 (subtract 3) 78 = 74x (divide by 74) 78/74 = x x = ?

OpenStudy (vuriffy):

This one is the 39/37 value I found graphing. 1/2 of the values in this situation, but equal the same decimal.

OpenStudy (vuriffy):

Can calculate x for me in this case?

OpenStudy (deercult):

Well, in that case, I got another repeating decimal, '1.054...'

OpenStudy (vuriffy):

Yes, correct.

OpenStudy (deercult):

Alright, then, I've got it written down. Awful tired of repeating decimals, tbh. Sooo, for th' second one.. 9-8x = -((2x + 3)/9) Yeh? Or just -(2x + 3)/9, I suppose?

OpenStudy (vuriffy):

9-8x = -((2x + 3)/9) is correct.

OpenStudy (vuriffy):

Work it out, and tell me what you get?

OpenStudy (deercult):

Alright, then... 9-8x = -2x - 3 / -9 eh? And goin' on, -81+72x = -2x - 3 -81 = -74x - 3 -78 = -74x Which just gets an even worse repeating decimal, '0.948717...'

OpenStudy (vuriffy):

Hmm, I got 1.20, not a repeating decimal.

OpenStudy (deercult):

I did somethin' wrong along the way then, since even (-78)/(-74) is '1.054...' accordin' to my calculator. Crud.

OpenStudy (vuriffy):

Do you mind if I do a proof instead this time?

OpenStudy (deercult):

Not at all.

OpenStudy (vuriffy):

Okay, the original is: 9l9-8xl= 2x + 3 My method was graphing, and I found the value where they intersected, so. Let's see if my value of 1.20 is correct for the x value.

OpenStudy (vuriffy):

9l9-8(1.2)l = 2(1.2) + 3 9l9-9.6l= 2.4 + 3 9l-.06l = 5.4 (mod it - absolute value) 9 * 0.6 = 5.4 5.4 = 5.4

OpenStudy (vuriffy):

Oops, 9l-0.6l = 5.4*****

OpenStudy (deercult):

I see!

OpenStudy (vuriffy):

So, this result in the value of 1.054 (repeating .054) and 1.20 to be the x values, you may also do a proof for the 1.054 and you will get 5.10.... = 5.10...... I already did all the work, so both do work in this case.

OpenStudy (vuriffy):

results*

OpenStudy (deercult):

Ah, nice! Thanks, then.

OpenStudy (vuriffy):

Of course, anytime. If you need any further explanation on anything, please ask.